# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@ 3 Local relaxation with fluctuations
At the basis of hydrodynamics is the idea of separation of scales: some observables vary slowly in space and time, while other vary much more quickly. By “integrating out” the fast observables, there remains a theory for slow observables, which are the local conserved densities admitted by the system. The Boltzmann Gibbs principle for hydrodynamics at the Euler scale says that linear fluctuations of observables project onto those of densities under such a fast relaxation [6]. Its nonlinear version, dubbed relaxation of fluctuations [25], gives access to large deviations. Beyond the Euler scale, it is well known that one must add fluctuations and diffusive effects. In this section we give heuristic arguments giving such a nonlinear fluctuating Boltzmann-Gibbs principle. We consider the ballistic scaling of space and time, and its corrections in $1/\ell$, but most of this discussion is relevant to other scalings as well, such as the diffusive scaling, see [35, 36, 21, 20, 23].
@@ 3.1 Nonlinear fluctuating Boltzmann-Gibbs principle
A mesoscopic region in one dimension is an interval of size $L$ much larger than microscopic lengths $\ell_{\rm micro}$ but much smaller than variation lengths $\ell$,
@@
We may consider $\ell_{\rm micro}$ to be finite (say, set to 1). We write
@@
This precise form of the fluid cell is not mandatory; one may for instance consider a weight function $w({\rm x})$ such that $|w({\rm x})|<a{\rm e}^{-b|x|}$ for some $a,b>0$ and $\int_{\mathbb{R}}{\rm d}{\rm x}\,w({\rm x})=1$, and define
@@
We argue in Sec. 3.2 that a nonlinear, fluctuating Boltzmann-Gibbs principle holds. That is, coarse-grained observables can be written solely in terms of fields representing coarse-grained conserved densities, which we denote $q_{i}(x,t)$, and emergent Gaussian white noise fields $\xi_{o}(x,t)$’s representing the “forgotten” microscopic degrees of freedom for the observable $o$. Both of these families of fields fluctuate, and we represent averages in the corresponding emergent fluctuating theory as $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$. We propose that one can make the following replacement, which has three parts, microcanonical, diffusive and stochastic (noise):
$\overline{o}(x,t)\ \mbox{in $\langle\cdots\rangle_{\ell}$}$
$\longrightarrow$ $o(x,t):=\underbrace{\mathsf{o}^{\rm reg}(\underline{q}(x,t))}_{\text{microcanonical}}\underbrace{-\frac{\operatorname{sgn}t}{2\ell}\sum_{i}\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))\partial_{x}q_{i}(x,t)}_{\text{diffusive}}+\underbrace{\frac{1}{\ell}\xi_{o}(x,t)}_{\text{noise}}\ \mbox{in $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$}$
with simply
$\overline{q_{i}}(x,t)\longrightarrow q_{i}(x,t)$ (19)
for conserved densities. Here,
$\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0,\quad\langle\!\langle\xi_{o}(x,t)\xi_{o^{\prime}}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}=\langle\!\langle\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}(x,t)\big{)}\rangle\!\rangle\,\delta(x-x^{\prime})\delta(t-t^{\prime})$ (20)
for some noise covariance $\hat{\mathfrak{L}}_{o,o^{\prime}}(\underline{q}(x,t))$, function of the fluctuating densities. $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))$ are the “bare” diffusion functions for the observable $o$ which accounts for fluctuation-dissipation effects from the noise and are irreversible (thus the $\operatorname{sgn}(t)$ factor), and $\mathsf{o}^{\rm reg}(\underline{q}(x,t))$ is a regularisation of the microcanonical average (12) in the fluid cell which account for non-fluctuating, non-diffusive corrections due to finite size of the cell. As we will see, the latter is crucial in order to make sense of singularities that arise because of short-range correlations, either as introduced by the initial state, or by the noise. The meaning of the various terms in (3.1) is discussed at more length in Sec. 3.2. These terms will be evaluated in Sec. 4 in terms of equilibrium quantities.
Eq. (3.1) is closely related to nonlinear fluctuating hydrodynamics [9], and a principle of the form (3.1) was proposed in [25, Sec 3.3] as an extension of the Ballistic Macroscopic Fluctuation Theory (BMFT) [25] to include diffusive effects. However in both cases, the constitutive functions $\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}(x,t)\big{)}$ and $\hat{\mathfrak{D}}_{o}^{~{}i}(\underline{q}(x,t))$ representing the corrections to the Euler scale are not evaluated, and the necessary regularisation $\mathsf{o}^{\rm reg}(\underline{q}(x,t))$ is not discussed. The “bare” noise in (3.1) was also studied in [37]. In systems that are purely diffusive, for current observables $o=j_{k}$ only the diffusive and stochastic part remain, as the microcanonical part is independent of $\underline{q}(x,t)$. This is the basis for the Macroscopic Fluctuation Theory (MFT) [8], a large-deviation theory for fluctuations on diffusive scales.
Recall that $\ell$ is the macroscopic scale: it is present both in $\overline{o}(x,t)$, via the ballistic scaling of space and time, and in the state $\langle\cdots\rangle_{\ell}$, its length scale for variations. The replacement (3.1) gives the $-\ell$ asymptotic expansion of macroscopic physical quantities evolved in time $t$. For instance, the theory gives the asymptotic (distributional) expansion of connected correlation functions: with (3.1) for $o_{k}(x_{k},t_{k})$, we have
@@
with
$\langle\!\langle o_{1}(x_{1},t_{1}),\ldots,o_{n}(x_{n},t_{n})\rangle\!\rangle^{\rm c}_{\ell}=\ell^{1-n}S_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ell^{-n}\delta S_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ldots$ (22)
See [25, 26] for the general principles of this expansion and the leading Euler-scale scaling. Similarly, for the large deviations and first subleading correction of macroscopic observables such as total currents, we have
$\log\langle{\rm e}^{\lambda\int_{0}^{\ell T}{\rm d}\mathrm{t}\,j_{k}(0,\mathrm{t})}\rangle_{\ell}=\log\langle\!\langle{\rm e}^{\lambda\ell\int_{0}^{T}{\rm d}t\,j_{k}(0,t)}\rangle\!\rangle_{\ell}+\mathcal{O}(\ell^{-1})=\ell F_{k}(\lambda,T)+\delta F_{l}(\lambda,T)+\mathcal{O}(\ell^{-1}).$ (23)
We leave for future work the evaluation of such correlation functions and large-deviation functions, and in particular the corrections to the Euler-scale projection formula developed in [26].
Crucially, the state $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$, which we have not specified except for (20), is in fact completely determined by the initial state correlations $S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$ and $\delta S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$, the replacement (3.1) (where all parts are specified as per Sec 4), and the conservation laws
@@
Indeed, this is because these then provide the dynamics and hence the correlations of conserved densities at later times: this is through the nonlinear fluctuating hydrodynamic equation
$\partial_{t}q_{i}(x,t)+\partial_{x}\Big{(}\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))-\frac{\operatorname{sgn}t}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~{}k}(\underline{q}(x,t))\partial_{x}q_{k}(x,t)+\frac{1}{\ell}\xi_{j_{i}}(x,t)\Big{)}=0$ (25)
where we use the more standard notation
$\hat{\mathfrak{D}}_{i}^{~{}k}:=\hat{\mathfrak{D}}_{j_{i}}^{~{}k}$ (26)
for the diffusion functions of current observable $j_{i}$.
One way of specifying the state more explicitly is using a path integral formulation such as in the MFT or the BMFT. For a principle of the form (3.1), this was proposed in [25, Sec 3.3], and the path integral formulation was written there, see in particular Eqs. (62)-(65) in that reference. Here, we will not make use of this path integral formulation and the associated large-deviation analysis – it turns out that for all large-deviation results and their subleading corrections, the techniques of cumulant expansion, initially suggested in [25], and fully developed at the Euler scale in [26], is sufficient.
@@
Usually one only considers noise terms associated to the local currents. However, in general, every observable $o$ has its associated, independent noise $\xi_{o}$, generically correlated to all other observables, including currents. It is simple to see from the evolution equation (25) that noise fields for observables that are not currents decouple from the theory, so may be omitted if one looks only at densities and currents.
@@
As usual, although $\xi_{o}$ are white noise fields, the theory is valid at the nonlinear level and for large deviations of local observables. Indeed, white noise fields occur because of the central limit theorem used in the argument of local relaxation to the microcanonical average, but the large-scale fluctuations of conserved densities give rise to nonlinearities and large deviations.

### v2

@@ 3 Local relaxation with fluctuations
At the basis of hydrodynamics is the idea of separation of scales: some observables vary and fluctuate slowly in space and time, while other do so much more quickly. By assuming that fast observables relax quickly, there remains a theory for slow observables, which are the local conserved densities admitted by the system. The Boltzmann Gibbs principle for hydrodynamics at the Euler scale says that linear fluctuations of observables project onto those of densities under such a fast relaxation [6]. Its nonlinear version says that every local observable is a fixed, non-fluctuating function of conserved densities, which are themselves fluctuating. This gives access to ballistic-scaling large deviations [7, 8, 43]. Beyond the Euler scale, one must add noise and diffusive effects. This is at the basis fluctuating hydrodynamics.
In this section we write the corresponding nonlinear fluctuating Boltzmann-Gibbs principle, and in order to justify its various parts, provide heuristic arguments for it. We consider the ballistic scaling of space and time (3), and its corrections in $1/\ell$. Most of this discussion is relevant to other scalings as well, such as the diffusive scaling, see [53, 54, 37, 36, 39].
@@ 3.1 Nonlinear fluctuating Boltzmann-Gibbs principle
A fluid cell is an abstract construct that is useful in order to determine the large-scale properties of many-body system. In one dimension, it is an interval of size $L$ much larger than microscopic lengths $\ell_{\rm micro}$ but much smaller than variation lengths $\ell$,
@@
Both $\ell_{\rm micro}$ (determined by the microscopic model) and $\ell$ (determined by the initial state) are measurable. $L$ is not, but it is a convenient theoretical concept. We may consider $\ell_{\rm micro}$ to be finite (say 1), by appropriate choice of microscopic units. We write
@@
This precise form of the fluid cell is not mandatory; one may for instance consider an even weight function $w({\rm x})=w(-{\rm x})$ such that $|w({\rm x})|<a{\rm e}^{-b|{\rm x}|}$ for some $a,b>0$ and $\int_{\mathbb{R}}{\rm d}{\rm x}\,w({\rm x})=1$, and define
@@
In order to simplify the form of the initial correlations in the state (12), it is important that the fluid-cell mean be balanced: the fluid-cell averaging is invariant under ${\rm y}\to-{\rm y}$.
We argue in Sec. 3.2 that a nonlinear, fluctuating Boltzmann-Gibbs principle holds. That is, coarse-grained observables can be written solely in terms of fluctuating fields representing coarse-grained conserved densities, which by a slight abuse of notation we denote $q_{i}(x,t)$, and emergent Gaussian white noise fields $\xi_{o}(x,t)$’s representing the “forgotten” fast microscopic degrees of freedom for the observable $o$. Both of these families of fields fluctuate, and we represent averages in the corresponding emergent fluctuating theory as $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$. The emergent fluctuating theory implicitly depends on $L$ via our choice of fluid cell, but it only gives “universal” results, independent of how $L$ is chosen, in the limit (15).
Before we express this emergent fluctuating theory, we mention that it is expected to describe the large-$\ell$ asymptotic expansion of all connected correlation functions at distinct macroscopic space-time coordinates $(x_{i},t_{i})\neq(x_{j},t_{j})\;\forall i\neq j$, as
@@
where $\langle\!\langle o_{1}(x_{1},t_{1}),\ldots,o_{n}(x_{n},t_{n})\rangle\!\rangle^{\rm c}_{\ell}$ are distributions, of the form
$\langle\!\langle o_{1}(x_{1},t_{1}),\ldots,o_{n}(x_{n},t_{n})\rangle\!\rangle^{\rm c}_{\ell}=\ell^{1-n}\mathsf{S}_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ell^{-n}\delta\mathsf{S}_{o_{1},\ldots,o_{n}}(x_{1},t_{1};\cdots;x_{n},t_{n})+\ldots.$ (20)
The leading large-deviation scaling $\ell^{1-n}$ is explained in [55, 8, 43].
The nonlinear Boltzmann-Gibbs principle says that the random variable $o(x,t)$ in the fluctuating theory $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$, representing the coarse-grained $\overline{o}(x,t)$, has three parts: microcanonical, diffusive and stochastic (Gaussian white noise) [^4] In correlation functions with $n\geq 3$, if the Euler hydrodynamics admits shocks, we expect (21) to hold almost everywhere in space-time. The specification “almost everywhere” is because at shocks the local relaxation arguments at its basis do not hold, and this, even at the diffusive order; and for $n\geq 3$, the nonlinear expansion required may be affected by shocks. In no-shock systems, such as with linear degeneracy, it is expected to hold everywhere except at colliding space-time positions. See the discussion in [43].. For all [^5] For $t<0$, the sign of the diffusive part is positive instead of negative, so in general we must replace $\longrightarrow-\operatorname{sgn}t$ in front of the diffusive term. $t>0$,
[…]
and in particular, for conserved densities,
$\overline{q_{i}}(x,t)\longrightarrow q_{i}(x,t).$ (22)
For the initial state (12), the noise $\xi_{o}(x,t)$ is exactly zero at $t=0$,
$\xi_{o}(x,0)=0,$ (23)
but it is nonzero generically for all $t>0$, see below. The initial state gives rise to the following one- and two-point correlations for the emergent conserved density fields,
$\langle\!\langle q_{i}(x,0)\rangle\!\rangle_{\ell}$ $=$ $\langle q_{i}\rangle_{\underline{\beta}(x)}+\mathcal{O}(\ell^{-2})$ (24)
$\langle\!\langle q_{i}(x,0)q_{j}(x^{\prime},0)\rangle\!\rangle_{\ell}$ $=$ $\ell^{-1}\mathsf{C}_{ij}(x)\delta(x-x^{\prime})+\mathcal{O}(\ell^{-3})$
where $\mathsf{C}_{ij}(x)$ is the covariance matrix evaluated within the state $\underline{\beta}(x)$. Note that $n$-point functions receive $\mathcal{O}(\ell^{-1})$ relative corrections, PT symmetry and because the fluid-cell mean is balanced, as we show in Appendix B.
The fluctuating theory $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$ is completely determined by the initial correlations $S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$ and $\delta S_{q_{i_{1}},\ldots,q_{i_{n}}}(x_{1},0;\cdots,x_{n},0)$, the replacement (21), and the conservation laws
@@
Indeed, these then provide the dynamics and hence the correlations of conserved densities at later times: this is through the nonlinear fluctuating hydrodynamic equation (here for $t>0$)
$\boxed{\partial_{t}q_{i}(x,t)+\partial_{x}\Big{(}\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))-\frac{1}{2\ell}\sum_{k}\hat{\mathfrak{D}}_{i}^{~k}(\underline{q}(x,t))\,\partial_{x}q_{k}(x,t)+\frac{1}{\ell}\xi_{i}(x,t)\Big{)}=0}$ (26)
where we use the simplified notation
$\hat{\mathfrak{D}}_{i}^{~k}:=\hat{\mathfrak{D}}_{j_{i}}^{~k},\quad\xi_{i}:=\xi_{j_{i}}.$ (27)
The replacement (21) along with (26) gives the large-$\ell$ asymptotic expansion (20).
Let us discuss the various elements in (21). The quantities $\hat{\mathfrak{D}}_{o}^{~i}(\underline{q})$ are the “bare” diffusion functions for the observable $o$, which account for fluctuation-dissipation effects from the noise, They represent irreversible effects (thus the $\operatorname{sgn}(t)$ factor in general). $\mathsf{o}^{\rm reg}(\underline{q})$ is the finite-size-corrected microcanonical average (11) in the fluid cell. It involves a “regularisation”, which is crucial in order to make sense of singularities that arise because of short-range correlations, as introduced by the initial state, Eq. (24).
Noise fields $\xi_{o}(x,t)$ are linear functions of $o$. As discussed in different contexts in [56, 57, 58, 59, 60], physically meaningful noise is to be treated in the Itô convention on time. That is, given $\mathcal{F}(t)=\{q_{i}(x^{\prime},t^{\prime}):x^{\prime}\in{\mathbb{R}},\,t^{\prime}\in[0,t],i\in\mathcal{I}\}$, the noise fields $\xi_{o}(x,t)$, for all observables $o$ (including currents $j_{i}$) and for all $x\in{\mathbb{R}}$, are such that $\xi_{o}(x,t){\rm d}t={\rm d}w_{o}(x,t):=w_{o}(x,t+{\rm d}t)-w_{0}(x,t)$ are centered Gaussian forward increment, with
$\langle\!\langle{\rm d}w_{o}(x,t){\rm d}w_{o^{\prime}}(x^{\prime},t)|\mathcal{F}(t)\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}(x,t)\big{)}\delta(x-x^{\prime}){\rm d}t$ (28)
for some symmetric noise covariance $\hat{\mathfrak{L}}_{o,o^{\prime}}(\underline{q})$, function of the densities.
Note that by this formula, the noise covariance $\hat{\mathfrak{L}}_{o,o^{\prime}}(\underline{q})$ is such that for any set of local observables $\mathcal{S}$, the matrix $\hat{\mathfrak{L}}(\underline{q})$ with elements $\hat{\mathfrak{L}}_{o,o^{\prime}}(\underline{q}):(o,o^{\prime})\in\mathcal{S}\times\mathcal{S}$ is non-negative. For $o=q_{i}$ a conserved density, according to (22) the noise must vanish, $\hat{\mathfrak{L}}_{q_{i},q_{i}}(\underline{q})=0$ (therefore $\hat{\mathfrak{L}}_{q_{i},o}(\underline{q})=0\,\forall\,o$ by the Cauchy-Schwartz inequality).
This gives the noise fields correlations at equal times. In order to express them at different times, we need to take away the explicit dependence of their covariance on the densities. Indeed, these are fluctuating, and correlate non-trivially with previous histories of noise fields because of the fluctuating hydrodynamic equation (26). For this purpose, fix a set of observables $\mathcal{S}$ not linearly related to each other and with non-vanishing noise. Then $\mathcal{S}$ is a basis for its span and the matrix $\hat{\mathfrak{L}}(\underline{q})$ is strictly positive on this basis. Define
$\hat{\xi}_{o}=\sum_{o^{\prime}\in\mathcal{S}}\Big{(}\sqrt{\hat{\mathfrak{L}}^{-1}}\Big{)}_{oo^{\prime}}\xi_{o^{\prime}}.$ (29)
Then the associated increments ${\rm d}\hat{w}_{o}(x,t):=\hat{\xi}_{o}(x,t){\rm d}t$ satisfy
$\langle\!\langle{\rm d}\hat{w}_{o}(x,t){\rm d}\hat{w}_{o^{\prime}}(x^{\prime},t)|\mathcal{F}(t)\rangle\!\rangle_{\ell}=\delta_{o,o^{\prime}}\delta(x-x^{\prime}){\rm d}t,\quad o,o^{\prime}\in\mathcal{S}.$ (30)
It is natural to assume that these are now independent, delta-correlated noise fields over all times,
$\langle\!\langle\hat{\xi}_{o}(x,t)\rangle\!\rangle_{\ell}=0,\quad\langle\!\langle\hat{\xi}_{o}(x,t)\hat{\xi}_{o^{\prime}}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}=\delta_{o,o^{\prime}}\delta(x-x^{\prime})\delta(t-t^{\prime}),\quad o,o^{\prime}\in\mathcal{S}.$ (31)
In general, correlations $\langle\!\langle\xi_{o}(x,t)\xi_{o^{\prime}}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}$ at different times will be very non-trivial, and $\xi_{o}(x,t)$ is non-Gaussian, with non-zero higher cumulants at different times. However, for the purpose of the fluctuating theory as a theory for the leading and first sub-leading order for correlation functions, Eqs. (19), (20), noise averages only occur in evaluating the diffusive $\mathcal{O}(\ell^{-1})$ relative correction. At this order, the cumulant expansion makes the noise covariance non-fluctuating, evaluated on the average densities. Hence, from this viewpoint $\xi_{o}(x,t)$ may be considered Gaussian, and we have
$\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0,\quad\langle\!\langle\xi_{o}(x,t)\xi_{o^{\prime}}(x^{\prime},t^{\prime})\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\langle\!\langle\underline{q}(x,t)\rangle\!\rangle_{\ell}\big{)}\,\delta(x-x^{\prime})\delta(t-t^{\prime}).$ (32)
When used in conjunction with the fluctuating hydrodynamic equation (26), this makes the latter not a standard stochastic PDE, but rather a self-consistent stochastic PDE, where the noise statistics is fixed by the averages densities solving this evolution equation in a self-consistent manner.
The meaning of the various terms in (21), as well as the emergence of the Itô convention, is discussed heuristically in Sec. 3.2. These terms will be evaluated in Sec. 4 by evaluating, with this fluctuating theory, particular equilibrium quantities including the Green-Kubo formula.
@@
Eq. (21) is the general form expected in fluctuating hydrodynamics (see e.g. [23]). However, the constitutive functions $\hat{\mathfrak{L}}_{o,o^{\prime}}\big{(}\underline{q}\big{)}$ and $\hat{\mathfrak{D}}_{o}^{~i}(\underline{q})$ representing the corrections to the Euler scale must be evaluated, as well as the regularisation $\mathsf{o}^{\rm reg}(\underline{q})$, which is typically not discussed. In systems that are purely diffusive, for current observables $o=j_{k}$ only the diffusive and stochastic part remain, as the microcanonical part is independent of $\underline{q}(x,t)$. This is the basis for the Macroscopic Fluctuation Theory (MFT) [11]. Without regularisation, diffusive and noise term, the result is the Ballistic Macroscopic Fluctuation Theory [8]. These are large-deviation theories, which avoid singularities due to the noise and initial conditions.
@@
Usually one only considers noise terms associated to the local currents, and takes $\mathcal{S}=\{j_{i}:i\in\mathcal{I}\}$. However, in general, every observable $o$ has its associated, independent noise $\xi_{o}$, generically correlated to all other observables, including currents. It is simple to see from the evolution equation (26) that noise fields for observables that are not currents decouple from the theory, so may be omitted if one looks only at densities and currents.